OpenStudy (anonymous):

sin^2x + sinx + cos^2x - 1

OpenStudy (imstuck):

What, exactly, are you doing with this?

OpenStudy (anonymous):

I have to simplify but I don't know how

OpenStudy (mathmale):

Please post the instructions for this problem. Note that (sin x)^2 is a better way to express "the square of the sine function" than is sin^2x. Rewrite this problem using (sin x)^2 and (cos x)^2, please.

OpenStudy (imstuck):

As far as I can tell, sin^2 x + cos^2 x = 1, so when you combine those into "1", you end up with 1 + 1 + sin(x) which is 2 + sin(x). That's what I can see as far as simplifying goes.

OpenStudy (mathmale):

I agree with that. The best way to approach the simplification of the given expression is to re-arrange the terms, so that you have (sin x)^2 + (cos x)^2 written in that manner (adjacent to each other), and can apply the identity presented by IMStuck.

OpenStudy (mathmale):

Mind writing the whole thing out, then rearranging the terms as indicated above, and then doing the final simplification?

OpenStudy (anonymous):

I think it should be 1-1+sin(x)= sin(x)

OpenStudy (anonymous):

oh I see!! so you have to cancel everything in parenthesis first and then finish

OpenStudy (imstuck):

nelsonjedi, I think you're right! I DO think it should it should be 1 - 1 + sin(x) = sin(x). You're right; I missed the -1 and had it as a +1.

OpenStudy (anonymous):

Been there done that

OpenStudy (anonymous):

i have couple more problems! can you help?

OpenStudy (mathmale):

Sure, so long as you'd please post them separately from this one.